OpenStudy (anonymous):

The stopping distance d of a car after the brakes are applied varies directly as the speed r. If a car traveling at 80 mph can stop in 360ft how many feet will it take the same car to stop when traveling at 70 mph?

OpenStudy (anonymous):

varies directly: as a increases, b increases. let a1 be 80mph and a2 be 70mph. Let b1 be 360ft and b2 be x. 80/360=70/x 80/360(70)=1/x 360(70)/80=x 315=x

OpenStudy (anonymous):

it's only a matter of ratio and proportion.

OpenStudy (anonymous):

The correct answer is 2205/8